OpenStudy (anonymous):

A total of 396 tickets were sold for the school play. They were either adult tickets or student tickets. There were 54 fewer student tickets sold than adult tickets. How many adult tickets were sold?

OpenStudy (p0sitr0n):

396 = x + x + 54 x : # of student tickets 396 = 2x +54 x = 171 x+54 = 225 171 student and 225 adult tickets sold